Grande03:

Drag the tiles to the correct boxes to complete the pairs. Not all tiles will be used. Match each equation of the piecewise function represented in the graph to its corresponding piece of the domain.

Grande03:

princemagic:

whats the main q?

HeartTricker:

ok so 1, Range \[\le x < 2\] 2. range \[\times = 2\] 3. \[2< x < 3\] 4. \[3<x \le 5\]

HeartTricker:

your welcome :)